how can i make the quadratic equation y=x^2+6x+5 into a word problem
Answer by ankor@dixie-net.com(22740)

y=x^2+6x+5 into a word problem
:
We have a rectangle, length is labeled (x+5), width is labeled (x+1)
Find the area (y)
y = (x+5)*(x+1)
Foil
y = x^2 + x + 5x + 5
y = x^2 + 6x + 5